OpenStudy (anonymous):

Reina and Sam are multiplying (84)*5 (73)*9.

OpenStudy (clalgee):

Reina and Sam's answer will be 275,40

OpenStudy (anonymous):

@Clalgee Reina's Work Sam's Work (84)5 (73)9 = 84 + 573 + 9 = 89712 (84)5 (73)9 = 84⋅573⋅9 = 820727

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

i think sam is correct